OpenStudy (peterg2001):

help will medal Which of the following is a solution of the equation 9 - 2x = -1? -5 5 4 -4

OpenStudy (peterg2001):

@shinebrightlikeadimon

OpenStudy (anonymous):

Substitute all the numbers for x, and see which equation is correct.

OpenStudy (anonymous):

9 - 2x = -1 -9 - 9 2x = -10 /2 /2 x = -5

OpenStudy (anonymous):

ill help :)

OpenStudy (anonymous):

or i wont haha.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x=-5 is wrong

OpenStudy (anonymous):

i meant 5

OpenStudy (anonymous):

simple mistake >.<

OpenStudy (anonymous):